55. For two factors-starchy or sugary, and green base leaf or white base leaf-the following counts for the progeny of self-fePlease only solve (a) and (b), thank you very much!
Answer #1

Solis e know that MLE of o for the given problem is c=o-036 Now using this MLE of o ceo and multiplying the resulting the prox’statistie - 03-23 270.3903 to.4768 to.257) - 9.0705 (6) %*-2) (005) = x60-05) = 10.6 (1-06 100% of X Z values are in this i